Determine if (3/4,-2) is a solution to -4x-7y=11
Answer by stanbon(75887)

Determine if (3/4,-2) is a solution to -4x-7y=11
---------------------
Substitute:
-4(3/4) -7(-2) = 11
----
-12 + 14 = 11
----
False
=====================
Cheers,
Stan H.
=====================